OpenStudy (anonymous):

What is the slope of the line through the points (5, 2) and (5, -3)?

OpenStudy (anonymous):

there is no slope, because the \(x\) value is the same in both coordinates

OpenStudy (anonymous):

When I plug in my numbers I got 0... Is that correct?

OpenStudy (anonymous):

it is "the change in \(y\) over the change in \(x\) , but since \(x\) does not change, the question makes no sense

OpenStudy (anonymous):

no that is not correct

OpenStudy (anonymous):

I don't understand then :(

OpenStudy (anonymous):

it is "undefined" because the zero is in the denominator

OpenStudy (anonymous):

So then how is the line like through these points?

OpenStudy (anonymous):

if you compute you get \[\frac{-3-2}{5-5}=\frac{-5}{0}\] but that is not a number, you cannot divide by zero, i.e. you cannot have a zero in the denominator

OpenStudy (anonymous):

it is a vertical line

OpenStudy (anonymous):



OpenStudy (anonymous):

a vertical line does not have a slope this is not the same as saying "the slope is zero" it means there is no slope for that line

OpenStudy (anonymous):

I see what you're saying now, that makes sense.

OpenStudy (anonymous):

okay good